1/2

John Howard Payne,

Home. sweet home

’Mid pleasures and palaces though we may roam, Be it ever so humble, there’s no place like home; A charm from the sky seems to hallow us there, Which, seek through the world, is ne’er met with elsewhere. Home, home, sweet, sweet home! There’s no place like home, oh, there’s no place like home!

An exile from home, splendor dazzles in vain; Oh, give me my lowly thatched cottage again! The birds singing gayly, that come at my call – Give me them – and the peace of mind, dearer than all! Home, home, sweet, sweet home! There’s no place like home, oh, there’s no place like home!

I gaze on the moon as I tread the drear wild, And feel that my mother now thinks of her child, As she looks on that moon from our own cottage door Thro’ the woodbine, whose fragrance shall cheer me no more. Home, home, sweet, sweet home! There’s no place like home, oh, there’s no place like home!

How sweet ’tis to sit ’neath a fond father’s smile, And the caress of a mother to soothe and beguile! Let others delight mid new pleasures to roam, But give me, oh, give me, the pleasures of home. Home, home, sweet, sweet home! There’s no place like home, oh, there’s no place like home!

To thee I’ll return, overburdened with care; The heart’s dearest solace will smile on me there; No more from that cottage again will I roam; Be it ever so humble, there’s no place like home. Home, home, sweet, sweet, home! There’s no place like home, oh, there’s no place like home!

…

2/2

Verkinto de tiu ĉi Angla poemo estas John Howard Payne (∗1791-06-09 – †1852-04-10).

Arg-1008-2038 (2014-05-16 14:16:47)

Pri la poeto vidu la retejon http: // en. wikipedia. org/ wiki/ John_Howard_Payne , pri la kanto http: // en. wikipedia. org/ wiki/ Home! _Sweet_Home! . La melodion verkis Henry Rowley Bishop (18 November 1786 – 30 April 1855). Pri tiu vidu la retejon http: // en. wikipedia. org/ wiki/ Henry_ Bishop_( composer) . Vidu ankaŭ: http: // www. ompersonal. com. ar/ ompoetry/ payne1. htm . La kanto povas esti aŭdata en http: // www. poemhunter. com/ poem/ home-sweet-home/ kaj en https: // www. youtube. com/ watch? v= TM0Sft7Kjic .

http://www.poezio.net/version?poem-id=1008&version-id=,2038